Incident
Seaforth Channel Incident

Incident Start Date
October 13, 2016 01:13

Current Date/Time
October 21, 2016 09:00

Location
Seaforth Channel, approximately 11nm west of Bella Bella

Situation
Tug Nathan E. Stewart and Barge DBL 55 ran hard aground on the rocks at 10/13/16 (01:13 PDT) yesterday morning. The unit was located in Canadian Waters at 52.14.46N, 128.23.160W at the entrance to Sea Forth Channel. The tug remains aground on Athlone Island, approximately 11.5 miles from Bella Bella by water. The barge was empty of cargo and has since been moved to a safe anchorage, at the mouth of Dundavan Inlet. Water ingress has been detected in the number 2,3, and 5 void tanks; all other barge voids and all cargo compartments have been inspected and there is no sign of water ingress. The tug is aground in shallow water and is partially submerged (only the mast is showing) in 9 metres of water. After the incident it was reported that the 1Center, and 1Starboard fuel tanks were breached. The tug was loaded with 59,924 gallons of diesel fuel. Kirby was able to confirm that 6554 gallons of diesel fuel were pumped from the tug into tank 6 on the barge before the pumps failed due to water in the engine room. A manifest of additional liquids aboard at 90% capacity has been prepared, consisting of the following: 639 gals lube oil, 550 gals hydraulic oil, 550 gals gear oil, 969 gals dirty bilge. Tug is considered in a stable position; it is sitting on the bottom with a roughly 12* list to port, and is not currently moving in the swell.

Priorities/Objectives
• Ensure the safety of citizens and response personnel
• Control the source
• Secure the vessel and refloat/remove from the marine environment
• Maximize protection of environmentally-sensitive areas, prioritizing of sample collection
• Contain and recover spilled material
• Remove oil from impacted areas
• Recover and rehabilitate injured wildlife
• Manage coordinated response effort
• Keep the public & stakeholders informed of response activities
• Clear and consistent communication to the community
• Minimize economic impacts from the spill to the community and the environment
• Minimize local business interruption

Unified Command
Jim Guidry (Responsible Party Incident Commander)
William Housty (Heiltsuk On Scene Commander)
Jeff Brady (Canadian Coast Guard Federal Incident Commander)
John Kervel (BC Provincial On Scene Commander)

Reports

SAFETY
• As of 06:50 hours, 21/10/16, all small and medium sized vessels have been given permission by Unified Command to resume work. They were stood down yesterday (20/10/16) at 12:40 hours due to safety concerns as a result of weather.
• Vessels have been instructed to make sure that any small boats launched check in with the BARTLETT to ensure they can be accounted for
• Handheld UHF radios have been distributed as needed, working with the repeater installed yesterday; ICP and on-water crews report good reception
WCMRC vessels are fitted with air monitoring resources and will report any safety concerns
• Contingency plans have been developed to shelter vessels in case of deteriorating weather
• Site Specific Safety Plan is being continuously developed and is to be followed by all responders
• Personal Protective Equipment stocks will be located on the MAR-ELL-MIST and the BARTLETT to ensure continued availability on scene; sorbent rolls will be available to limit contamination of vessel decks and to reduce slipping hazard
• A Decon Plan has been developed, and a vessel decontamination area is available at the GULF RIVAL

FISHERIES CLOSURE
•Fisheries and Oceans Canada have put in place a fisheries closure for the area as of Friday
•For more information: Elysha Gordon 250-756-7192

INCIDENT COMMAND POST
• ICP is operating out of the Heiltsuk Nation Tribal Council chambers during Day and Night Operations

STAGING AREA
• Staging Area has been established at Shearwater

RESPONSE
20/10/16, 06:30 hours – 20/10/16, 18:30 hours
• Divers worked within the tug today (20/10/16) and have tapped the 1Centre tank; divers located product trapped against the ceiling of the engine room, and vacuumed this product out before pumping the 1Centre tank; the remaining tanks will be prioritized for tapping and pumping in order of tank volume in the following days.
• NASP overflight at 10:00 hours today (20/10/16) observed an estimated 23 litres (6 US gallons) of product on the water
• Sheen was observed outside of the secondary containment boom.
• It is recognized that there is uncertainty in NASP volume estimates.
• Overflight assets include NASP aircraft, CCG helicopter and Kirby helicopter
o Overflight tasking will be coordinated through the ICP, and will be focusing on tactical operations support, observations, and supporting SCAT surveys
o Accountability flight to review the operating area was conducted and will be conducted end of day going forward as necessary
• Replacement and collection of sorbents has been ongoing; crews have collected 237 bags of solid waste so far; bags are being stored in double lined storage bins for eventual transport to Vancouver for disposal

20/10/16, 18:30 hours – 21/10/16, 06:30 hours
• All vessels were stood down or stood by on scene for the night between 20/10/16, 18:30 and 21/10/16, 06:30.
• At 01:13 hours on 21/10/16, the Central Coaster, standing-by on scene, reported a section of secondary containment boom that had been set onto a rock formation during low tide. The Eagle Bay deployed the FLIR to evaluate the situation. The FLIR did not indicate any product outside of the secondary containment. The GM Penman was alerted of the situation and put on standby, prepared to launch at any point from Bella Bella. Drones were launched at 03:45 with a delay in time due to weather conditions. The drone FLIR did detect coloration, potentially indicating sheen outside the secondary boom barrier. As a result, the Eagle Bay was deployed to actively skim in the area down current from where the boom had lifted off the rocks. The Eagle Bay did not detect any product in their skimmers. The GM Penman remained on stand-by as to not impede operations for today (21/10/16, 06:30 – 18:30).

21/10/16, 06:30 hours – 21/10/16, 18:30 hours
• Today (21/10/16) weather depending, existing general purpose boom will remain in place and be surrounded by 1500’ of Kepner boom. An additional layer of general purpose boom will be added. Sorbent boom, lining the inside of the general purpose boom, will be replaced and monitored every hour to ensure booms integrity and effectiveness.
• Dive teams will be deploying from the docks at 07:00 hours on 21/10/16. They are expected to be on scene by 08:30 hours. Their first priority is to survey the boat and assess its condition. After which vacuuming product inside the vessel will commence.
• Operations crews will continue to maintain existing booming strategies. The condition of protection boom will be assessed and replaced as required. Boom will continue to be maintained and tended throughout the day, recovering and replacing soiled sorbents and operating skimming resources as required.
• Unified Command has agreed that no protection booming strategies are to be removed until the topic has been discussed and the removal approved
• Skimming vessels, Eagle Bay and GM Penman will remain on scene and active with 24/7 shifts.
• Contingency plans have been developed by Operations in order to immediately respond to any release during pumping operations, both during the day and at night; response and surveillance assets will be on scene, and all other available resources will be on standby for redirection to the scene if necessary in case of a release
• NASP overflights will occur again today at 10:00 hours and 14:00 hours, weather depending.

SHORELINE CLEANUP ASSESSMENT TECHNIQUES
20/10/16, 06:30 hours – 20/10/16, 18:30 hours
• Yesterday (20/10/16), pre-SCAT surveys occurred around the mouth of Gale Channel. SCAT teams had and will continue to have representation from Heiltsuk Nation, BC Parks, CCG, BC Ministry of Environment and Polaris.
• SCAT crews were stood down at 12:40 hours (20/10/16) due to weather.
• Areas surveyed near the incident site on the south side of Seaforth Channel did show signs of stranded oil. As a result, shoreline crews have been working in the area, following Shoreline Treatment Recommendations approved by Unified Command.

20/10/16, 18:30 hours – 21/10/16, 06:30 hours
• SCAT crews were stood down over night and will resume work at 07:00 hours on the 21st of October.

21/10/16, 06:30 hours – 21/10/16, 18:30 hours
• Pre-SCAT surveys will occur at the following locations: south shoreline of Seaforth Sound, working west along Athlone Island. Crews will continue to work according to the SCAT Plan, which remains under review by Unified Command.
• Treated areas along the south side of Seaforth Channel will be inspected by SCAT teams today (21/10/16).
• Surveys being conducted for the time being are known as ‘pre-SCAT’ surveys; SCAT surveys cannot start until all potential oil has been removed from the water and the tug

WILDLIFE
20/10/16, 06:30 hours – 20/10/16, 18:30 hours
• Wildlife teams conducted wildlife monitoring on the south shore of Seaforth Channel, west of the spill release site. The team then moved to the north side of Seaforth Channel and surveyed the south, east and north shores of Ivory Island and surrounding waters, including Paul Anchorage. The team then moved to the southwest side of Spiller Channel. Monitoring operations in Spiller Channel were limited in scope due to a midday weather stand down (12:40 PM, 20/10/16).
• Nine sea otters were observed in Paul Anchorage (north of Ivory Island). One whale was reported seen north of the casualty site, on the north side of Seaforth Channel. NO VISIBLY IMPACTED OR OILED WILDLIFE HAS BEEN OBSERVED THROUGHOUT THE AREA.

20/10/16, 18:30 hours – 21/10/16, 06:30 hours
• Wildlife crews were stood done overnight.

21/10/16, 06:30 hours – 21/10/16, 18:30 hours
• Wildlife crews will resume surveys today (21/10/16).
• Wildlife teams – weather depending – will survey the southwest side of Spiller Channel and the south shore of Seaforth Channel.
• Wildlife Hotline has been activated: 877-323-3985
• Operations crews have been instructed to report any wildlife sightings through MAR-ELL-MIST for relay to the ICP.
• Wildlife crews will continue to follow-up with Heiltsuk Guardian Watchmen who have been monitoring the area for oiled wildlife and have observed some wildlife in the spill area over the last few days.
• Public and responders are again urged not to attempt to handle any oiled wildlife and follow proper wildlife handling and notifications protocols.

ENVIRONMENT UNIT
• Sampling crews have resumed operations today (21/10/16); Sampling crews were stood down at 1240 yesterday (20/10/16) due to weather
• BC Ministry of Environment is cooperating with Heiltsuk Nation and being supported remotely by Environment Canada
• Environment Canada has been supporting contingency planning for hot tapping and pumping operations with trajectory modelling in case of a release
• BC Ministry of Environment is participating with Heiltsuk and Kirby to implement the Sampling Plan; Sampling Plan is hoped to be submitted for approval by Unified Command today
• BC Ministry of Environment is developing a Waste Plan to support Operations

MEDIA
• Spill information website has been activated
www.spillresponsebc.ca

Seaforth Channel Site Specific Forecast

Forecast Issue Time: 06:00 AM PDT 10/21/2016

seaforth-0600

Seaforth Channel Specific Synoptic Discussion

The forecast as of 06:00 hours on the 21st of October is the following: An approaching trough of low pressure will produce showers and thunderstorms with strong gusty winds today. The showers will be heavy at times through today and tonight, causing reduced visibilities. Saturday will see improving weather and diminishing winds as the trough weakens and moves inland.

Seaforth Channel Weather: On scene update at 08:00 AM 21/10/16
Current Conditions:
Wind: Southeasterly 9-10 knots
Precipitation: Medium Drizzle
Sea State: Long period swell approximately 0.5-1.0m in height from the west
Visibility: 4-6NM

Marine Weather Statement

Issued 3:57 AM PDT 21 October 2016
A 990 millibar low pressure centre over Northwestern Bowie this morning will move northward towards the Alaskan Panhandle today. The low will bring strong to gale force southerlies over most northern waters this morning. Southwest gales will veer to westerly gales today over the northern offshore waters as the low moves by to the north. Winds will ease to moderate to strong tonight.

Central Coast Marine Forecast (Environment Canada)

WINDS

Issued 04:00 AM PDT, 21 October 2016
GALE WARNING IN EFFECT. Wind southerly 35-45 knots diminishing to southeast 25-35 this morning and to the southeast 15-25 early this afternoon. Wind increasing to southeast 25-35 late this evening then diminishing to southeast 15-25 overnight. Wind becoming southeast 20-30 Saturday (22/10/16).

WAVES

Issued at 04:00 AM PDT, 21 October 2016
Today (Friday 21/10/16), Tonight and Saturday – Seas 3 to 5 metres subsiding to 2-3 after midnight.

WEATHER & VISIBILITY

Issued 04:00 AM PDT, 21 October 2016
Periods of rain. Risk of thunderstorms this morning (21/10/16). Visibility 1 mile or less in rain.

EXTENDED FORECAST

Issued 04:00 AM PDT 21 October 2016
SUNDAY (23/10/16) Wind southeast 35-45 knots diminishing to southeast 15-25 in the afternoon.
MONDAY (23/10/16) Wind southeast 25-35 knots.
TUESDAY (24/10/16) Wind southeast 15 to 25 knots.

GALE WARNING IN EFFECT:
Central Coast from McInnes Island to Pine Island
Issued 04:00 AM PDT 21 October 2016

 

TIDES
tides-0600